A garden hose can fill a swimming pool in 3 days, and a larger hose can fill the same pool in 19 hours. How long will it take to fill the empty pool if both hoses are used together for the same amount of time?
Answer by Cromlix(4381)

Garden hose can fill the pool in 3 days (72 hours)
Large garden hose can fill the pool in 19 hours
1/72 + 1/19 = 1/x
Common denominator = 1368x
19x + 72x = 1368
91x = 1368
x = 15hours and 2 minutes
Hope this helps.
:-)